# Addressing Positivity Violations in Continuous Interventions through Data-Adaptive Strategies

### v2

@@ 2.3.2 Diagnostic measure: the non-overlap ratio
To assess positivity in practice, we introduce the non-overlap ratio. This diagnostic summarizes, at the population level, the extent to which a candidate intervention value $a$ falls outside the conditional support region defined by the HDRs. Formally, we define a function $\tau:\mathcal{A}\to[0,1]$ by
$\tau(a):=\int\mathbbm{1}\{a\notin\mathcal{A}_{\alpha}(\mathbf{L})\}\,dP_{\mathbf{L}}.$ (5)
Here, $\tau(a)$ represents the proportion of individuals whose covariate profiles $\mathbf{L}$ render the intervention value $a$ unsupported, i.e., lying outside the practical support region where the positivity assumption is satisfied.
The non-overlap ratio takes values between 0 and 1:
@@
  • Values near 1 indicate that $a$ is largely unsupported, reflecting severe positivity violations and suggesting that causal inferences at this value are unreliable.
@@ 3.1 Standard Causal Dose–Response Curve
Estimand 1 (Standard) The standard causal dose–response curve is the function $m^{\text{standard}}:\mathcal{A}\to\mathcal{Y}$ defined by $m^{\text{standard}}(a)=\mathbb{E}(Y^{a}),$ which can be identified as $m^{\text{standard}}(a)$ $\overset{(i)}{=}\,\mathbb{E}\!\left[\mathbb{E}(Y^{a}\mid A=a,\mathbf{L})\right]$ (6) $\overset{(ii)}{=}\,\mathbb{E}\!\left[\mathbb{E}(Y\mid A=a,\mathbf{L})\right]$ $=\int\mathbb{E}(Y\mid A=a,\mathbf{L}=\mathbf{l})\,dP_{\mathbf{L}}(\mathbf{l}),$ where equality (i) follows by the law of iterated expectations and conditional exchangeability, and equality (ii) by consistency and the positivity assumption (Equation 1).
The standard causal dose–response curve, reliant on the positivity assumption in Equation 1, faces serious challenges due to frequent violations, particularly with continuous interventions. As the intervention range widens, regions of the confounder space often have limited or no support, so that $\hat{f}(a\mid\mathbf{L})$ approaches zero. In such cases, identification fails and estimation becomes unstable.
This difficulty is magnified in practice because continuous exposures span infinitely many potential values, making it impossible to observe all treatment levels across every confounder pattern in finite samples. Consequently, estimation procedures are forced to extrapolate into sparsely or never observed regions of the intervention space. Such extrapolation produces unstable estimates, inflates standard errors, and introduces bias, thereby compromising both the validity and the generalizability of the estimated dose–response relationship.
@@
3.2 Existing Estimand to Address Positivity Violations:
Weighted Causal Dose–Response Curve
@@ 3.2 Existing Estimand to Address Positivity Violations:
Estimand 2 (Weighted CDRC) Schomaker et al. [31] propose the weighted CDRC, defined as $m^{\text{weighted}}(a)$= $\,\mathbb{E}\!\left(Y^{a}\,\middle|\,f(a\mid\mathbf{L})>c\right)P\!\big(f(a\mid\mathbf{L})>c\big)$ (7) $+\mathbb{E}\!\left(Y^{a}\,w(a\mid\mathbf{L})\,\middle|\,f(a\mid\mathbf{L})\leq c\right)P\!\big(f(a\mid\mathbf{L})\leq c\big),\qquad\forall a\in\mathcal{A},$ which admits the identification formula $m^{\text{weighted}}(a)=\int\mathbb{E}(Y\mid A=a,\mathbf{L}=\mathbf{l})\,w(a,\mathbf{l})\,dP_{\mathbf{L}}(\mathbf{l}),$ (8) where the weight function $w(a,\mathbf{l})$ is given by $w(a,\mathbf{l})=\begin{cases}1,&\text{if }f(a\mid\mathbf{l})>c,\\[6.0pt] \dfrac{f(a\mid\mathbf{l})}{f(a)},&\text{if }f(a\mid\mathbf{l})\leq c.\end{cases}$ (9)
The weighted CDRC mitigates positivity violations by coinciding with the standard estimand in regions with adequate support, and by redefining it in regions with limited support through a projection based on conditional densities. This ensures identifiability and avoids extrapolation into unsupported regions. However, it changes the target parameter: the resulting curve no longer represents the causal effect of assigning all individuals to treatment level $a$, but rather a blend of causal and weighted associational components. As a consequence, the weighted CDRC may diverge from the scientific estimand of interest, and its causal interpretation becomes less transparent.
More broadly, this approach highlights a fundamental trade-off in addressing positivity violations: one can either preserve the original scientific estimand but accept estimation that is potentially unstable, highly variable, or biased in regions with sparse data, or redefine the estimand to guarantee identifiability at the cost of causal interpretability. The weighted CDRC takes the latter path, producing well-defined estimates that avoid extrapolation into unsupported regions, but in doing so it targets a modified scientific question. In applied settings, researchers must therefore balance the benefits of improved precision and robustness against the loss of a direct causal interpretation when deciding whether to adopt the weighted CDRC.
@@
3.3 Novel Estimands to Address Positivity Violations: Causal Dose–Response Curve with Data-Adaptive Intervention
To address these challenges, we propose a novel estimand that mitigates the impact of positivity violations by focusing on well-supported regions of the intervention space, guided by the diagnostic tool proposed in the previous section. This approach relaxes the strict positivity requirement by focusing on empirically supported regions of the intervention space, while maintaining interpretability of the resulting causal estimand.
3.3.1 Data-Adaptive Intervention
Unlike static interventions, which assign a fixed treatment $a$ to all units, we introduce a more refined approach called the data-adaptive intervention. This strategy customizes the intervention for each individual based on their confounders and the observed data.
For a given oracle intervention value $a$, the idea can be summarized schematically as
[…]
Formally, let $\mathcal{A}_{\alpha}(\mathbf{l};P)$ denote the highest density regions (HDRs) of treatment values supported by the distribution $P$ given covariates $\mathbf{l}$, as defined in Section 2.2.
We define the data-adaptive intervention operator
$\mathop{}\!\mathrm{d}:\mathcal{A}\times\mathcal{L}\times\mathcal{P}\rightarrow\mathcal{A},$
where $\mathcal{P}$ denotes the space of candidate distributions. The operator assigns to each $(a,\mathbf{l},P)$ a feasible treatment value via
$\mathop{}\!\mathrm{d}(a,\mathbf{l};P)=\begin{cases}a,&\text{if }a\in\mathcal{A}_{\alpha}(\mathbf{l};P),\\[6.0pt] h(a,\mathbf{l};P),&\text{otherwise},\end{cases}$ (11)
where $h:\mathcal{A}\times\mathcal{L}\times\mathcal{P}\to\mathcal{A}$ is a substitution function satisfying
$h(a,\mathbf{l};P)\in\mathcal{A}_{\alpha}(\mathbf{l};P)\quad\text{whenever }a\notin\mathcal{A}_{\alpha}(\mathbf{l};P).$ (12)
The corresponding counterfactual outcome under the data-adaptive intervention is denoted
$Y^{\mathop{}\!\mathrm{d}}\equiv Y^{\mathop{}\!\mathrm{d}(a,\mathbf{l};P)}.$
In practice, the true distribution $P$ is unknown and replaced by the empirical distribution $P_{n}$, which makes the rule inherently data-adaptive. For notational convenience, write
$\mathcal{A}_{\alpha}(\mathbf{l}):=\mathcal{A}_{\alpha}(\mathbf{l};P_{n})\quad\text{and}\quad h(a,\mathbf{l}):=h(a,\mathbf{l};P_{n}).$
Then the data-adaptive intervention rule becomes
$\mathop{}\!\mathrm{d}(a,\mathbf{l};P_{n})=\begin{cases}a,&\text{if }a\in\mathcal{A}_{\alpha}(\mathbf{l}),\\[6.0pt] h(a,\mathbf{l}),&\text{otherwise}.\end{cases}$ (13)
3.3.2 The Most Feasible Intervention
A natural choice for the substitution function $h$ is to reassign unsupported interventions to the nearest feasible value, ensuring minimal deviation from the intended intervention:
$h(a,\mathbf{l})=\underset{a^{*}\in\mathcal{A}_{\alpha}(\mathbf{l})}{\arg\min}\,|a^{*}-a|.$ (14)
@@ 3.3.2 The Most Feasible Intervention
Figure 2: Illustration of the most feasible intervention defined in Equations 13 and 14. Panel (a) shows the conditional density of treatment values given covariates, with HDRs (green) representing well-supported regions and grey areas denoting poorly supported regions. Target interventions (red arrows) are either retained if they fall inside HDRs or reassigned to the nearest feasible value within HDRs (green arrows). Panel (b) extends this principle across confounder strata, where targets (red circles) are reassigned to the closest supported value (green circles) whenever they fall outside HDRs, while feasible targets remain unchanged. Together, the panels illustrate the principle of the most feasible intervention: retain the target when supported, and otherwise make the smallest necessary adjustment to ensure feasibility.
3.3.3 Proposal 2: A Novel Estimand Based on Feasible Interventions
Estimand 3 (Feasible) With the definitions of support and the substitution function in Equations 4 and 14, the proposed estimand is $m^{\text{feasible}}(a)=\mathbb{E}\!\left(Y^{\mathop{}\!\mathrm{d}}\right),$ (15) where $\mathop{}\!\mathrm{d}$ is defined in Equation 13 and $h$ in Equation 14. Equivalently, $m^{\text{feasible}}(a)=\mathbb{E}\!\Big[Y^{a}\mathbbm{1}\{a\in\mathcal{A}_{\alpha}(\mathbf{L})\}+Y^{\,h(a,\mathbf{L})}\mathbbm{1}\{a\notin\mathcal{A}_{\alpha}(\mathbf{L})\}\Big].$ (16)
In words, each unit either (i) receives the target intervention $a$ if $a$ lies in its supported region $\mathcal{A}_{\alpha}(\mathbf{L})$, or (ii) is reassigned to the nearest feasible value $h(a,\mathbf{L})$ otherwise. The resulting curve averages these counterfactuals across the population, yielding a causal dose–response curve that closely aligns with the original estimand while remaining identifiable in the presence of positivity violations. By limiting substitutions to the minimal adjustment necessary, this approach preserves interpretability and avoids the distortions introduced by extrapolation into unsupported regions.
@@ 3.3.3 Proposal 2: A Novel Estimand Based on Feasible Interventions
Both the weighted CDRC and the feasible CDRC aim to address positivity violations by restricting attention to regions with sufficient empirical support. However, they adapt the estimand in fundamentally different ways. The weighted CDRC of Schomaker et al. [31] applies a fixed cutoff $c$ to the conditional density and redefines the target through projection weights. This ensures identifiability without requiring extrapolation, but it alters the causal parameter: in regions of support, the curve coincides with the standard CDRC, while in unsupported strata it relies on weighted associational components. As Schomaker et al. emphasize, this means the weighted curve sticks to the causal question where feasible but falls back on associations where necessary, yielding a compromise between interpretability and stability. By contrast, the feasible CDRC employs a unit-specific, data-adaptive threshold and modifies intervention assignments explicitly by reassigning unsupported values to the closest feasible ones. This produces a transparent intervention rule targeting the well-defined quantity $E(Y^{\mathop{}\!\mathrm{d}})$. In short, the weighted CDRC trades some causal interpretability for broader identifiability, whereas the feasible CDRC preserves interpretability while adapting directly to the observed support.
@@
Figure 2b illustrates how the most feasible intervention strategy operates at the individual level, conditional on confounders. The vertical axis represents confounder values, and the horizontal axis represents the intervention. The shaded green band marks the highest density regions, where intervention values are well supported by the data, while the surrounding grey area corresponds to lowest density regions with little or no support. The red vertical line indicates the oracle intervention target, which may not be feasible for all individuals. Units whose oracle intervention lies outside the highest density regions (shown as red circles) cannot be assigned this target directly. Instead, the proposed rule reassigns them to the nearest feasible value within the highest density regions, shown by the green circles and arrows. By contrast, units whose oracle target already falls within the highest density regions (green circles on the red line) retain their original assignment. Black squares represent the observed data (intervention–confounder pairs), from which the highest density regions are estimated. Overall, the figure demonstrates that the most feasible intervention rule systematically reassigns unsupported targets to the closest feasible values, ensuring that every individual receives an intervention assignment compatible with the observed data distribution.
@@
The proposed CDRC with the most feasible intervention addresses these issues by adapting the intervention target to stay within regions where data are sufficiently abundant and well-supported. Instead of forcing an unrealistic target that fails to align with real-world conditions, this approach adjusts the intervention based on what is practically achievable and observable within the data. This has two key benefits: first, it reduces bias caused by data sparsity and positivity violations, and second, it ensures that the causal effects being estimated are relevant and actionable in practice.

### v3

@@ 3.3.2 Diagnostic measure: the non-overlap ratio
To assess positivity in practice, we introduce the non-overlap ratio. This diagnostic summarizes, at the population level, the extent to which a candidate intervention value $a$ falls outside the conditional support region defined by the HDRs. For a fixed $\alpha\in(0,1)$, define $\tau_{\alpha}:\mathcal{A}\to[0,1]$ by
$\tau_{\alpha}(a):=\int_{\mathcal{L}}\mathbbm{1}\{a\notin\mathcal{A}_{\alpha}(\mathbf{l})\}\,dP_{\mathbf{L}}(\mathbf{l})\;=\;\mathbb{P}\big(a\notin\mathcal{A}_{\alpha}(\mathbf{L})\big).$ (5)
Here, $\tau_{\alpha}(a)$ represents the proportion of individuals whose covariate profiles $\mathbf{L}$ render the intervention value $a$ unsupported, i.e., lying outside the practical support region where positivity is regarded as plausible at level $\alpha$.
The non-overlap ratio takes values in $[0,1]$:
@@
  • Values near 1 indicate that $a$ is largely unsupported, reflecting severe positivity violations.
@@ 4 Estimands to Address Positivity Violations
Throughout this section, we view each estimand as a functional of an observed-data law $P$ for $O=(\mathbf{L},A,Y)$, and denote its value under the true law $P_{0}$. Specifically, for an estimand $m^{\dagger}$ ($\dagger\in\{\text{standard},\text{trim},\text{feasible}\}$), we write $m^{\dagger}(P)$ for the functional and $m^{\dagger}_{0}:=m^{\dagger}(P_{0})$ for the target under $P_{0}$. Expectations and densities under $P$ are denoted $\mathbb{E}_{P}$ and $f_{P}$, with $\mathbb{E}_{0}$ and $f_{0}$ corresponding to $P_{0}$.
@@ 4.1 Standard Causal Dose–Response Curve
Estimand 1 (CDRC) The standard causal dose–response curve is $m^{\text{standard}}_{0}:\mathcal{A}\to\mathcal{Y}$ defined by $m^{\text{standard}}_{0}(a)=\mathbb{E}_{0}\!\left[Y^{a}\right],\qquad a\in\mathcal{A}.$ Identification of $m^{\text{standard}}_{0}(a)$ relies on the following standard conditions for each $a\in\mathcal{A}$: 1. Consistency: $Y=Y^{a}$ when $A=a$. 2. Conditional exchangeability: $Y^{a}\perp A\mid\mathbf{L}$. 3. Positivity: $f_{0,A\mid\mathbf{L}}(a\mid\mathbf{l})>0$ for $P_{0,\mathbf{L}}$-almost every $\mathbf{l}$. Under these conditions, $m^{\text{standard}}_{0}(a)=\mathbb{E}_{0}\!\left[\mathbb{E}_{0}\!\left(Y\mid A=a,\mathbf{L}\right)\right].$ (6)
The standard causal dose–response curve is conceptually appealing, but its identification and stable estimation hinge on the positivity assumption in the box above. In many applied continuous-intervention settings, this condition can be plausible for some intervention values and implausible for others. When $f_{0,A\mid\mathbf{L}}(a\mid\mathbf{l})$ is near zero in parts of the covariate space, estimation of $m^{\text{standard}}_{0}(a)$ becomes fragile in the sense that estimation must rely on stronger modeling extrapolations across $\mathbf{l}$, and common estimators can exhibit high variance and sensitivity to model misspecification because they are effectively learning about $Y$ at $(A=a,\mathbf{L}=\mathbf{l})$ from sparse data.
Importantly, these difficulties are not a consequence of continuity alone, but of limited empirical support for specific target intervention values within covariate strata. This is well illustrated by the CHAPAS-3 motivation (Section 2): EFV concentrations are downstream of dosing and pharmacokinetics and are therefore only partially controllable. As a result, some concentration levels are rarely observed for certain metabolic and clinical profiles, and the empirical support for EFVt is concentrated in a mid-range, with substantially weaker support at the extremes (Section 6; Figure 6a). In such regions, estimating the standard concentration–response curve at extreme $a$ values is expected to be more dependent on extrapolation and correspondingly less stable, motivating the diagnostics and alternative estimands that explicitly restrict attention to well-supported interventions.
@@
4.2 Existing Estimand to Address Positivity Violations: Weighted Causal Dose–Response Curve
@@
Estimand 2 (Weighted CDRC) Schomaker et al. [31] propose the weighted CDRC, which is defined as $m^{\text{weighted}}_{0}(a)$= $\,\mathbb{E}_{0}\!\left(Y^{a}\,\middle|\,f_{0}(a\mid\mathbf{L})>c\right)P_{0}\!\big(f_{0}(a\mid\mathbf{L})>c\big)$ (7) $+\mathbb{E}_{0}\!\left(Y^{a}\,w_{0}(a\mid\mathbf{L})\,\middle|\,f_{0}(a\mid\mathbf{L})\leq c\right)P_{0}\!\big(f_{0}(a\mid\mathbf{L})\leq c\big),\qquad\forall a\in\mathcal{A},$ where the weight function is $w_{0}(a,\mathbf{l})=\begin{cases}1,&\text{if }f_{0}(a\mid\mathbf{l})>c,\\[6.0pt] \dfrac{f_{0}(a\mid\mathbf{l})}{f_{0}(a)},&\text{if }f_{0}(a\mid\mathbf{l})\leq c.\end{cases}$ (8) Under consistency and conditional exchangeability (as in Estimand 1), and assuming $f_{0}(a)>0$ (so that $w_{0}(a,\mathbf{l})$ is well defined), $m^{\text{weighted}}_{0}(a)$ admits the identification formula $m^{\text{weighted}}_{0}(a)=\mathbb{E}_{0}\!\Big(\mathbb{E}_{0}\!\big(Y\mid A=a,\mathbf{L}\big)\,w_{0}(a,\mathbf{L})\Big).$ (9)
The weighted CDRC can be viewed as a modification of the standard CDRC that leaves the target unchanged on the well-supported subset $\{f_{0}(a\mid\mathbf{L})>c\}$, while altering the contribution of covariate strata with low conditional density by introducing the weight $w_{0}(a,\mathbf{L})$. This yields an estimand that is defined for all $a\in\mathcal{A}$ and is less sensitive to near-violations of positivity at treatment levels $a$ that are rarely observed within some covariate strata.
At the same time, the resulting target differs from the standard intervention estimand $m^{\text{standard}}_{0}(a)=\mathbb{E}_{0}(Y^{a})$ whenever $P_{0}\{f_{0}(a\mid\mathbf{L})\leq c\}>0$. In those regions, $m^{\text{weighted}}_{0}(a)$ averages the conditional mean $\mathbb{E}_{0}(Y\mid A=a,\mathbf{L})$ using weights that depend on the treatment density, so it no longer corresponds to the counterfactual mean under the single intervention that deterministically sets $A$ to $a$ for everyone. Rather, it is a density-weighted functional that coincides with the standard CDRC only where $f_{0}(a\mid\mathbf{L})>c$. Consequently, the interpretation of the weighted curve depends on the choice of $c$ and on how one views the density-based reweighting in the low-support region.
More broadly, estimand modifications designed to address limited overlap trade direct correspondence to the standard “set $A$ to $a$ for everyone” target for improved stability when empirical support is sparse. The weighted CDRC is one principled example of this strategy: it provides a well-defined target across the full range of $a$, but its scientific meaning differs from $m^{\text{standard}}_{0}(a)$ whenever the modification is active. In applied work, it is therefore useful to report the weighted CDRC alongside overlap diagnostics and to be explicit about the implied target when $f_{0}(a\mid\mathbf{L})$ falls below the cutoff.
@@
4.3 Novel Estimands to Address Positivity Violations: Feasible Dose–Response Curve (FDRC)
To mitigate the impact of positivity violations, we propose a causal estimand that restricts interventions to covariate-specific regions of the treatment space that are well supported under the data-generating mechanism. The key idea is to replace a target intervention level $a$ by a covariate-adaptive feasible value whenever $a$ falls outside the supported region for a given covariate profile, thereby reducing reliance on extrapolation while retaining a population-level causal target.
$d_{0,\alpha}(a,\mathbf{l})=\begin{cases}a,&\text{if }a\in\mathcal{A}_{\alpha}(\mathbf{l};P_{0}),\\[6.0pt] \Pi_{\mathcal{A}_{\alpha}(\mathbf{l};P_{0})}(a),&\text{if }a\notin\mathcal{A}_{\alpha}(\mathbf{l};P_{0}),\end{cases}$ (10)
where $\Pi_{\mathcal{A}_{\alpha}(\mathbf{l};P_{0})}(a)$ is the closest supported value to $a$:
$\Pi_{\mathcal{A}_{\alpha}(\mathbf{l};P_{0})}(a)=\arg\min_{u\in\mathcal{A}_{\alpha}(\mathbf{l};P_{0})}|u-a|.$ (11)
Assume that for $P_{0}$-almost every $\mathbf{l}$, the set $\mathcal{A}_{\alpha}(\mathbf{l};P_{0})$ is nonempty and closed (for example, when $\mathcal{A}$ is compact). Then the minimum in (11) exists, and if multiple values are equally close, we apply a fixed tie-breaking rule (e.g., choose the smallest minimizer); see Appendix LABEL:apd:proof_fdrc:remarks. Intuitively, this intervention rule keeps the target dose $a$ when it is supported for covariate profile $\mathbf{l}$, and otherwise replaces it with the nearest supported dose. For each fixed $a$, the map $\mathbf{l}\mapsto d_{0,\alpha}(a,\mathbf{l})$ is therefore an individualized deterministic intervention rule.
@@ 4.3 Novel Estimands to Address Positivity Violations: Feasible Dose–Response Curve (FDRC)
Figure 2: Illustration of the most feasible intervention defined in Equations (10) and (11). Panel (a) shows the conditional density of treatment values given covariates, with HDRs (green) representing well-supported regions and grey areas denoting poorly supported regions. Target interventions (red arrows) are either retained if they fall inside HDRs or reassigned to the nearest feasible value within HDRs (green arrows). Panel (b) extends this principle across confounder strata, where targets (red circles) are reassigned to the closest supported value (green circles) whenever they fall outside HDRs, while feasible targets remain unchanged. Together, the panels illustrate the principle of the most feasible intervention: retain the target when supported, and otherwise make the smallest necessary adjustment to ensure feasibility.
Estimand 3 (Feasible Dose-Response Curve) Fix $\alpha\in(0,1)$. For each target level $a\in\mathcal{A}$, let $d_{0,\alpha}(a,\mathbf{l})$ be the most feasible intervention operator defined in (10)–(11), so that the intervention assigns $A\leftarrow d_{0,\alpha}(a,\mathbf{L})$ and $d_{0,\alpha}(a,\mathbf{l})\in\mathcal{A}_{\alpha}(\mathbf{l};P_{0})$ for $P_{0,\mathbf{L}}$-almost every $\mathbf{l}$. Let $Y^{d_{0,\alpha}(a,\cdot)}$ denote the corresponding counterfactual outcome. The proposed feasible dose–response curve is $m_{0}^{\text{feasible}}(a)=\mathbb{E}_{0}\!\left\{Y^{d_{0,\alpha}(a,\cdot)}\right\},\qquad a\in\mathcal{A}.$ (12) Identification assumptions. For each $a\in\mathcal{A}$, assume: 1) Consistency: $Y=Y^{a}$ when $A=a$ for all $a\in\mathcal{A}$; in particular, $Y=Y^{d_{0,\alpha}(a,\cdot)}$ when $A=d_{0,\alpha}(a,\mathbf{L})$. 2) Conditional exchangeability: $Y^{a}\perp A\mid\mathbf{L}$ for all $a\in\mathcal{A}$. 3) Well-definedness: for $P_{0,\mathbf{L}}$-almost every $\mathbf{l}$, $\mathcal{A}_{\alpha}(\mathbf{l};P_{0})$ is nonempty (and closed), so that the projection in (11) exists; if the minimizer is not unique, ties are resolved by a deterministic measurable rule so that $(a,\mathbf{l})\mapsto d_{0,\alpha}(a,\mathbf{l})$ is measurable. Under 1)–3) [^1] A formal derivation of this identification result is provided in Appendix D.5. , $m_{0}^{\text{feasible}}(a)=\mathbb{E}_{0}\!\left[\mathbb{E}_{0}\!\left\{Y\mid A=d_{0,\alpha}(a,\mathbf{L}),\,\mathbf{L}\right\}\right].$ Unlike the standard CDRC, identification does not require positivity at the target level $a$ for all covariate strata; instead, by construction $d_{0,\alpha}(a,\mathbf{l})\in\mathcal{A}_{\alpha}(\mathbf{l};P_{0})$, so the estimand only intervenes on treatment levels that are well supported within each $\mathbf{l}$-stratum.
Relation to dynamic interventions.
The map $(a,\mathbf{l})\mapsto d_{\alpha}(a,\mathbf{l})$ is an individualized (covariate-dependent) intervention operator. For each fixed $a$, the induced map $\mathbf{l}\mapsto d_{\alpha}(a,\mathbf{l})$ can be viewed as a (single-time) dynamic treatment rule. A key distinction from the dynamic regimes most commonly studied in causal inference is that our rule is defined through a covariate-specific supported set $\mathcal{A}_{\alpha}(\mathbf{l};P_{0})$, which is a functional of the conditional exposure distribution under $P_{0}$. Thus the covariate dependence is introduced for feasibility: for each $(a,\mathbf{l})$, the rule coincides with the static intervention “set $A$ to aa” when $a\in\mathcal{A}_{\alpha}(\mathbf{l};P_{0})$, and otherwise maps $a$ to a nearby supported value.
@@
Both the weighted CDRC and the FDRC aim to address positivity violations by restricting attention to regions with sufficient support. However, they adapt the estimand in fundamentally different ways. The weighted CDRC of Schomaker et al. [31] applies a fixed cutoff $c$ to the conditional density and redefines the target through projection weights. This ensures identifiability without extra assumptions, but it alters the causal parameter: in regions of support, the curve coincides with the standard CDRC, while in unsupported strata it relies on weighted associational components. As Schomaker et al. emphasize, this means the weighted curve sticks to the causal question where feasible but falls back on associations where necessary, yielding a compromise between interpretability and stability. By contrast, the FDRC employs a unit-specific threshold and modifies intervention assignments explicitly by reassigning unsupported values to the closest feasible ones. This produces a transparent intervention rule targeting the well-defined quantity $E(Y^{d_{0,\alpha}(a,\cdot)})$. In short, the weighted CDRC trades some causal interpretability for broader identifiability, whereas the FDRC preserves interpretability while adapting directly to the actual support.
@@
Figure 2b illustrates how the most feasible intervention strategy operates at the individual level, conditional on confounders. The vertical axis represents confounder values, and the horizontal axis represents the intervention. The shaded green band marks the highest density regions, where intervention values are well supported by the data, while the surrounding grey area corresponds to lowest density regions with little or no support. The red vertical line indicates the oracle intervention target, which may not be feasible for all individuals. Units whose oracle intervention lies outside the highest density regions (shown as red circles) cannot be assigned this target directly. Instead, the proposed rule reassigns them to the nearest feasible value within the highest density regions, shown by the green circles and arrows. By contrast, units whose oracle target already falls within the highest density regions (green circles on the red line) retain their original assignment. Black squares represent the observed data (intervention–confounder pairs), from which the highest density regions are estimated. Overall, the figure demonstrates that the most feasible intervention rule systematically reassigns unsupported targets to the closest feasible values, ensuring that every individual receives an intervention assignment compatible with the data distribution.
@@
In this setting, the feasible dose–response curve (FDRC) answers a simple, real-world question: for each concentration level we aim for, what outcome would we expect on average if children who can realistically reach that level are set to it, while children for whom that level is not realistically attainable are instead moved to the closest concentration they can realistically attain? In other words, the FDRC describes what would happen if we target a given concentration but allow individual “fallback” adjustments when the target is out of reach. It therefore matches the usual dose–response curve where targets are broadly attainable, and it differs mainly at extreme targets that are unrealistic for some children.
@@
The proposed feasible dose–response curve addresses these issues by adapting the intervention target to remain within regions with adequate empirical support. Instead of imposing a target that is weakly supported by the observed data, this approach modifies the intervention according to what is practically achievable and observable in the data. This has two key benefits: first, it can improve robustness and stability of estimation in the presence of sparsity and practical positivity violations (relative to unsupported targets), and second, it ensures that the causal effects being estimated are relevant and actionable in practice.